News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 13 May 2022 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

A real-time procedure was executed on May 7 to evaluate the telemetry link margin using the transmitter in low power mode when Chandra is near orbit apogee. This activity was in support of a planned operational configuration change to protect DSN resources during the upcoming epoch of low perigee.

Chandra passed through the 10th and 11th eclipses of the season on May 8 and 11 respectively, with nominal power and thermal performance. A real-time procedure was executed on May 8 to dump and clear the EPS glitch counters following the eclipse.

A Chandra press release was issued on May 12 describing the coordination of multiple telescopes, including Chandra, which observed the Milky Way's giant black hole simultaneously with the Event Horizon Telescope (EHT). This combined effort gave insight into what is happening farther out than the field-of-view of the EHT. These data will help astronomers better understand the complex process of "accretion" where material falls towards and into the black hole. X-rays from Chandra reveal hot gas that was blown away through winds from the black hole known as Sagittarius A*. For details see:

The schedule of targets for the next week is shown below and includes an observation of SDSS J150921.68+030452.7 coordinated with HET.

        Radiation Belts                     May 16
        DELSJ092347.12+0402    ACIS-S
        GW170817               ACIS-S       May 17
        DELSJ110421.59+2134    ACIS-S
        SDSSJ150921.68+0304    ACIS-S
        GDH17                  ACIS-I
        GDH17                  ACIS-I       May 18
        Abell2744              ACIS-I
        GDH17                  ACIS-I
        GDH17                  ACIS-I
        PSZ2G018.32-28.50      ACIS-I
        Radiation Belts                     May 19
        JWST-DTDF-1            ACIS-I
        hetej1900.1-2455       ACIS-S
        NGC3081                ACIS-S       May 20
        2RXSJ001122.0-21143    ACIS-S
        2FHLJ1745.1-3035       ACIS-I
        4U0114+65              ACIS-S/HETG
        GDH17                  ACIS-I       May 21
        KeplerSNR              ACIS-S/HETG
        Radiation Belts
        rhoOphA                ACIS-S/HETG
        SDSSJ012530.85-10273   ACIS-S       May 22
        ASASSN-14ko            ACIS-S
        SDSSJ095031.63+43290   ACIS-S
        Abell2029              ACIS-I
        UGC4393                ACIS-S
        Abell2029              ACIS-I
        [DAG2016]PSZ1-ARCG     ACIS-S       May 23


All spacecraft subsystems except HRC continued to support nominal operations.

Return to Status Reports